a) Find the slope of the line through the points (3, -7) and (-2, 5).
b) Write in point-slope form, Y - Y₁= M ( X - X₁), the equation of the line through the two points given in part a.
c) Convert the equation that you wrote in part b to slope-intercept form, Y = MX + B.

Answer by Cromlix(4381)

Hi there,
a) Gradient of line between (3,-7) and (-2,5)
Gradient = y(2) - y(1)/x(2) - x(1)
Gradient = 5 - (-7)/-2 - 3
Gradient = 12/-5
b) Using (-2,5)
Y - Y₁= M ( X - X₁)
Y - 5 = -12/5(X -(-2))
Y - 5 = -12/5(X + 4)
c) Y = -12/5xX - 48/5 + 25/5 (5 = 25/5)
Y = -12/5X - 23/5
You can multiply thro by 5 if you want whole numbers
5Y = -12X - 23
Hope this helps:-)